A 2.50- mu F capacitor is charged to 747 V and 6.8
Answer #1

here,

capaciatnce 1 , C1 = 2.5 * 10^-6 F and V1 = 747 V

capacitance 2 , C2 = 6.8 * 10^-6 F and V2 = 564 V

charge on 1 , Q1 = C1 * V1

Q1= 1.8675 * 10^-3 C

Q2 = C2 * V2 = 3.8352 * 10^-3 C

A)

the charge across each capacitor , Q = (Q1 + Q2)/2

Q = 2.85 * 10^-3 C

the potential difference across 1, V1 = Q/C1

V1 = 1141 V

B)

the potential difference across 2 , V2 = Q/C2

V2 = 419 V